If diameter of human eye is 2.5cm, find the power range of the eye
1
Expert's answer
2020-05-04T12:40:09-0400

For clear vision, the image must be on the retina. Let "d=2.5\,cm=0.025\,m", "d_0=\infty,\, d_1=0.25\,m" for diameter of human eye, distant vision and close vision respectively. We have "P_i=\frac{1}{d_i}+\frac{1}{d}" for power of a lens.


Hence the power of the eye ranges from "P_0=40\, D" (for distant totally relaxed vision) to "P_1=44\,D" (for close fully accommodated vision).